# Decoupled Data Consistency with Diffusion Purification for Image Restoration

### v5

@@ 4.1 Faster Diffusion Purification
Lastly, notice that this approximation may seem similar to CMs, as they both necessitate only a single NFE. However, CMs avoid some of the tradeoffs discussed previously, as CMs were designed specifically for single-step sampling. Nevertheless, we illustrate the effectiveness of our framework on both techniques, demonstrating that one can use any model that is more readily available.
4.2 Adaptation to Latent Diffusion
Latent diffusion models (LDMs) [24, 25] operate in a lower-dimensional latent space by training a pair of encoder and decoder [42] to address the slow inference time of diffusion models. Specifically, given an encoder $\mathcal{E}:\mathbb{R}^{d}\to\mathbb{R}^{r}$ and a decoder $\mathcal{D}:\mathbb{R}^{r}\to\mathbb{R}^{d}$ where $r\ll d$, LDMs reduce the computational complexity by mapping each $\bm{x}\in\mathbb{R}^{d}$ to $\bm{z}=\mathcal{E}(\bm{x})\in\mathbb{R}^{r}$ and performing the entire diffusion process in the latent space. Then, upon running the reverse steps within the latent space, one can pass the sampled latent code $\bm{z}_{0}$ through the decoder to obtain a synthesized image $\bm{x}_{0}=\mathcal{D}(\bm{z}_{0})$.
Unlike previous methods [20, 21], the flexibility of our framework allows us to (almost) seamlessly incorporate LDMs. Similar to Section 3.1, we alternate between enforcing data consistency and diffusion purification but both in the latent space:
$\widehat{\bm{v}}_{k}$ $=\texttt{DPUR}(\bm{z}_{k},\bm{s}_{\bm{\phi}},T_{k}),$ (17)
@@ 4.2 Adaptation to Latent Diffusion
Approach I: Enforcing DC in the Latent Space.
One way of enforcing data consistency (DC) is strictly in the latent space. Specfically, we can consider the modified problem
$\bm{z}_{k}$ $=\operatorname*{\text{arg~{}min}}_{\bm{z}}\frac{1}{2}\|\bm{y}-\mathcal{A}(\mathcal{D}(\bm{z}))\|_{2}^{2}+\frac{\mu}{2}\|\bm{z}-\widehat{\bm{v}}_{k-1}\|_{2}^{2},$
where $\widehat{\bm{v}}_{k-1}$ is the outcome of the previous diffusion purification step in the latent space. We remark that this problem is similar to the CSGM problem considered in [4].
Then, similar to Section 3.2, we can approximately solve the above problem by taking a limited number of gradient based steps on $\frac{1}{2}\|\bm{y}-\mathcal{A}(\mathcal{D}(\bm{z}))\|_{2}^{2}$ w.r.t. $\bm{z}$ starting from $\widehat{\bm{v}}_{k-1}$:
Before the diffusion purification step, we run an extra step on the solution obtained from Equation 18
$\bm{z}_{k}$ $\leftarrow\mathcal{E}(\mathcal{D}(\bm{z}_{k})),$ (19)
which we label “Re-Encode”. The rationale behind this seemingly unnecessary step is that due to the nonlinearity of the encoder $\mathcal{E}$, there exist infinitely many solutions $\bm{z}$ for the optimization problem in Equation 18. However, since the LDMs are trained only on $\bm{z}=\mathcal{E}(x)$, we need to first map $\bm{z}_{k}$ back to the image space with $\bm{x}_{k}=\mathcal{D}(\bm{z}_{k})$ and then re-encode $x_{k}$ as $z_{k}=\mathcal{E}(\bm{x}_{k})$.
Approach II: Enforcing DC in the Pixel Space.
Alternatively, we can also directly enforce the data consistency in the pixel space. Specifically, given $\widehat{\bm{v}}_{k-1}$ as the outcome of the previous diffusion purification step in the latent space, we can run
$\bm{x}_{k}$ $=\operatorname*{\text{arg~{}min}}_{\bm{x}}\frac{1}{2}\|\bm{y}-\mathcal{A}(\bm{x})\|_{2}^{2}+\frac{\mu}{2}\|\bm{x}-\mathcal{D}(\widehat{\bm{v}}_{k-1})\|_{2}^{2}.$
Again, we can approximately solve the above problem by taking a limited number of gradient based steps on $\frac{1}{2}||\bm{y}-\mathcal{A}(\bm{x})||_{2}^{2}$ w.r.t. $\bm{x}$ starting from $\mathcal{D}(\widehat{\bm{v}}_{k-1})$:
[…]
which is the same as Equation 10 with $\bm{v}_{k-1}=\mathcal{D}(\widehat{\bm{v}}_{k-1})$. Followed by this, we run
$\bm{z}_{k}$ $=\mathcal{E}(\bm{x}_{k}),$
@@
to use the encoder $\mathcal{E}$ to map $\bm{x}_{k}$ into the latent space $\bm{z}_{k}$ for the subsequent diffusion purification step in Equation 17.
@@
Compared to Approach I, it is worth noting that Approach II is much more efficient. This is because here we do not require back propogation through the deep decoder $\mathcal{D}$ for computing the gradient. In practice, both methods are effective for a range of image restoration problems. However, we observe that Approach II tends to be more suitable for specific types of problems like motion deblurring and Gaussian deblurring. This is because it has a better capability to produce $\bm{x}_{k}$’s with superior quality compared to Approach I. See Figures 15 and 14 in Section A.1 for more detail.
@@ 5 Experiments
In this section, we demonstrate the effectiveness of our method under various settings. First, we introduce the basic experiment setup in Section 5.1. Second, we show our main experimental results in Section 5.2. Finally, we provide an ablation study in Section 5.3 in terms of noise robustness, computational efficiency, and convergence.
@@ 5.1 Experimental Setup
Types of Image Restoration Problems.
To demonstrate the effectiveness of our algorithm, we consider the following image restoration problems:
  • 1)
    Image inpainting with a $100\times 100$ size box mask.
  • 2)
    Image super-resolution with 4×4$\times$4 × down sampling.
  • 3)
    Gaussian image deblurring with Gaussian kernel of size $61\times 61$ and standard deviation $\sigma=3.0$.
  • 4)
    Motion image deblurring with motion deblurring kernel size $61\times 61$ and intensity value $0.5$.
These image restoration problems are commonly explored in numerous relevant works [20, 21, 16, 13] and also arise in many practical applications.
Evaluation Datasets and Metrics.
To demonstrate the effectiveness of our method, we conducted experiments on the FFHQ [43] and LSUN-Bedroom [44] datasets, which have an image size of $256\times 256\times 3$. For CMs, we test our method only on the LSUN-Bedroom dataset as a proof-of-concept. Generally, images of bedrooms have more complicated structures than those in “face” image datasets, making them a good indicator of the efficacy of our algorithm.
To quantitatively evaluate the performance of the algorithms, we use standard metrics: (i) Peak Signal-to-Noise-Ratio (PSNR), (ii) Structural Similarity Index Measure (SSIM) [45] and (iii) Learned Perceptual Image Patch Similarity (LPIPS) [46]. These metrics are commonly used in existing works that evaluate image restoration techniques [21], and are shown to be an accurate indicator of reconstruction performance.
Figure 4: Pixel space diffusion models for solving various image restoration problems. These results demonstrate the effectiveness of our algorithm. We can consistently recover images with the highest PSNR.
@@
Ours Pixel (Version 1) $\bm{0.215}\pm 0.05$ $\underline{26.10}\pm 3.12$ $\underline{0.768}\pm 0.08$ $\underline{0.098}\pm 0.01$ $20.79\pm 2.66$ $\underline{0.880}\pm 0.02$
Ours Pixel (Version 2) $0.229\pm 0.05$ $\bm{26.92}\pm 3.33$ $\bm{0.800}\pm 0.07$ $\bm{0.094}\pm 0.02$ $\bm{23.66}\pm 3.42$ $\bm{0.906}\pm 0.02$
@@
Ours LDMs (Version 1) $\underline{0.244}\pm 0.05$ $\underline{26.52}\pm 3.21$ $\underline{0.778}\pm 0.08$ $\underline{0.215}\pm 0.04$ $\underline{23.75}\pm 3.32$ $\bm{0.831}\pm 0.05$
Ours LDMs (Version 2) $0.259\pm 0.05$ $\bm{26.88}\pm 3.28$ $\bm{0.789}\pm 0.08$ $0.239\pm 0.04$ $\bm{24.22}\pm 3.65$ $\underline{0.828}\pm 0.05$
Ours CMs $0.208\pm 0.05$ $26.17\pm 3.19$ $0.774\pm 0.08$ $0.108\pm 0.02$ $19.66\pm 2.48$ $0.874\pm 0.02$
Table 1: Quantitative results of super resolution and inpainting on the LSUN-Bedrooms dataset for both pixel-based and LDMs. The results colored in blue and in orange are for LDMs and CMs respectively, while the others are for pixel space diffusion models. The best results are in bold and second best results are underlined for pixel space models and LDMs respectively.
Comparison Baselines.
Since we conduct experiments involving both pixel space and latent space, we consider the following baselines:
  • Diffusion Models in Pixel Space. For the pixel space diffusion models, we compare our algorithm to DPS [16], Manifold Constrained Gradients (MCG) [13] and Plug-and-Play ADMM [47] (PnP-ADMM).
  • Diffusion Models in Latent Space. For the experiments with LDMs, we compare to Latent-DPS [21], ReSample [21], and Posterior Sampling with Latent Diffusion Models (PSLD) [20].
To ensure a fair comparison, we use pre-trained pixel-space diffusion models from Chung et al.[16] and Dhariwal et al.[10], as well as pre-trained CMs from Song et al. [19], for both the baseline techniques and our methods in the pixel space. Similarly, for diffusion models in the latent space, we use models provided by Rombach et al. [24], which were previously used in existing methods [21].
@@
Ours Pixel (Version 1) $\bm{0.196}\pm 0.05$ $\underline{27.13}\pm 3.26$ $\underline{0.804}\pm 0.07$ $\underline{0.065}\pm 0.01$ $\underline{33.56}\pm 3.92$ $\underline{0.947}\pm 0.02$
Ours Pixel (Version 2) $\underline{0.212}\pm 0.05$ $\bm{27.74}\pm 3.34$ $\bm{0.825}\pm 0.07$ $\bm{0.067}\pm 0.02$ $\bm{34.47}\pm 4.07$ $\bm{0.956}\pm 0.02$
@@
Ours LDMs (Version 1) $0.246\pm 0.05$ $\underline{26.08}\pm 3.02$ $\underline{0.766}\pm 0.07$ $\underline{0.145}\pm 0.03$ $\underline{28.87}\pm 3.77$ $\underline{0.856}\pm 0.06$
Ours LDMs (Version 2) $\bm{0.217}\pm 0.05$ $\bm{27.11}\pm 3.26$ $\bm{0.807}\pm 0.07$ $\bm{0.121}\pm 0.03$ $\bm{29.40}\pm 3.81$ $\bm{0.875}\pm 0.05$
Ours CMs $0.202\pm 0.05$ $26.91\pm 3.11$ $0.805\pm 0.07$ $0.061\pm 0.02$ $34.10\pm 3.80$ $0.953\pm 0.02$
Table 2: Quantitative results of Gaussian and motion deblurring on the LSUN-Bedrooms dataset for both pixel-based and LDMs. The results colored in blue and orange are for LDMs and CMs respectively, while the others are for pixel space diffusion models. The Best Results are in bold and second best results are underlined for pixel space models and LDMs respectively.
5.2 Main Experimental Results
Here, we present both quantitative and qualitative results for our proposed algorithm and the baselines. The algorithms are tested with clean measurements generated from $100$ randomly selected images from each of the FFHQ and LSUN-Bedrooms datasets. For conciseness, we only display results on the LSUN-Bedroom dataset and defer the results on FFHQ to Appendix B. We also defer the choices of hyperparameters in Section A.2.
5.2.1 Pixel Space Diffusion Models
In this section, we present the performance of our proposed method when using diffusion models trained in the pixel space. Recall that we have two versions of our method for the pixel space: vanilla reverse SDE sampling as well as an approximation method via Tweedie’s formula. We label the vanilla reverse sampling method as Version 1 (V1) and the Tweedie’s fomula approximation method as Version 2 (V2). We also demonstrate the use case of our algorithm using CMs as discussed in Section 4.1.
Figure 5: Demonstrating the differences between Versions 1 and 2 on super resolution (4×4$\times$4 ×). The two columns show the reconstructions with PSNRs and their corresponding “canny edges”, respectively.
The qualitative results are reported in Figure 4 and the quantitative comparison of pixel space diffusion model algorithms is shown in Table 1 and Table 2. In these tables, it is evident that our methods noted as V1 and V2 consistently outperform the existing baselines. However, even though the quantitative results imply that V2 is able to achieve higher performance on standard metrics such as PSNR and SSIM, the reconstructions are often blurrier and contain less perceptual detail compared to V1. As shown in Figure 5, the reconstruction of V1 has more high frequency information, visualized with “canny edges”, despite having lower PSNR. This phenomenon is commonly referred to as the “perception-distortion tradeoff” [48, 49].
Surprisingly, we observe that MCG seems to have a lot of variability in the results on the LSUN-Bedroom dataset, but performs well on FFHQ (shown in Appendix B). This might suggest that datasets such as faces are generally easier than datasets with more complicated structures such as bedrooms.
For the results with CMs, we present the quantitative results in Table 1 and  Table 2 highlighted in orange, with visual results in Figure 6. This highlights another use case of our algorithm, as it demonstrates that our algorithm can flexibly utilize CMs compared to existing methods.

### v6

@@ 3.4 DCDP with Latent Diffusion and Consistency Models
Approach II. Data Consistency in Pixel Space
Alternatively, we can directly enforce the data consistency in the pixel space. Specifically, given $\widehat{\bm{v}}_{k-1}$ as the outcome of the previous diffusion purification step, we can optimize the following objective:
$\bm{x}_{k}$ $=\operatorname*{\text{arg~{}min}}_{\bm{x}}\frac{1}{2}||\bm{y}-\mathcal{A}(\bm{x})||_{2}^{2}+\mu\|\bm{x}-\mathcal{D}(\widehat{\bm{v}}_{k-1})\|_{2}^{2}.$ (29)
Similar to that in the pixel space, (29) can be approximately solved by taking a limited number of gradient-based steps on $\frac{1}{2}||\bm{y}-\mathcal{A}(\bm{x})||_{2}^{2}$ when initialized from $\mathcal{D}(\widehat{\bm{v}}_{k-1})$, with
@@
which is equivalent to (18) with $\bm{v}_{k-1}=\mathcal{D}(\widehat{\bm{v}}_{k-1})$. Followed by this, we run $\bm{z}_{k}=\mathcal{E}(\bm{x}_{k})$ to map $\bm{x}_{k}$ into the latent space for the subsequent diffusion purification step in (25).
@@
Compared to Approach I, it is worth noting that Approach-II is much more efficient since it does not require back-propagation through the deep decoder $\mathcal{D}$ when computing the gradient. In practice, both methods are effective for a range of inverse problems. However, we find that Approach II is better suited for tasks like motion deblurring and Gaussian deblurring, as it consistently yields higher-quality reconstructions than Approach I (See Figures 16 and 15 in Appendix A).
DCDP with Consistency Models. Recently, Song et al. [17] proposed a class of diffusion models termed consistency models (CMs), which can effectively generate high-quality samples in a single step. CMs can be seamlessly incorporated into our framework. Specifically, with a pre-trained consistency model $\bm{f}_{\bm{\theta}}$, which maps a noisy sample $\mathbf{x}_{k}$ at time $t_{k}$ directly to its denoised estimate, the diffusion purification step in (20) can be performed in a single step as
$\bm{v}_{k}=\texttt{DPUR}(\bm{f}_{\bm{\theta}},\bm{x}_{k},t_{k}):=\bm{f}_{\bm{\theta}}(\bm{x}_{k},t_{k}).$ (31)
Performing diffusion purification with CMs may seem similar to the approximation via Tweedie’s formula, as they both require only a single NFE. However, CMs alleviate the blurriness issue brought by the Tweedie’s formula discussed in Section 3.3. This is because unlike the Tweedie’s formula that estimates the posterior mean, CMs are designed specifically for sampling from the data distribution with a single step. We illustrate the effectiveness of our framework on both techniques in Section 4.2, demonstrating that one can use any model that is readily available.
3.5 Comparison with Related Works
Before demonstrating the effectiveness of our approach in Section 4, we first discuss how it relates to recent closely related works. For a more comprehensive overview of other diffusion model-based methods, such as [38], we refer readers to the recent survey by Daras et al. [39] and to Section 2 of Zheng et al. [40].
Plug-and-Play Diffusion Priors. Our work, together with several concurrent studies [41, 42, 43], falls under the category of plug-and-play diffusion-based inverse problem solvers. These methods alternate between a data fidelity step and a prior step during inference. While both our approach and [41, 42] use implicit priors of diffusion denoisers, the latter directly performs denoising on $\bm{x}_{k}$ by assuming $\bm{x}_{k}$ contains a certain amount of pseudo noise, which might be a strong assumption in practice. Our method explicitly injects noise to $\bm{x}_{k}$ before performing denoising with either iterative sampling or one-step using Tweedie’s formula.
This explicit noise-then-denoise process—termed diffusion purification in our work—is also implicitly adopted in DAPS [43]. The main distinction lies in the data fidelity step: DAPS uses Langevin dynamics for posterior sampling, whereas our formulation in (17) can be viewed as a maximum a posterior (MAP) estimation. Hence, the two algorithms can be seen as complementary. As they were developed concurrently, we leave a thorough comparison of the two approaches for future work.
Lastly, the study presented by Luther et al. [44] shares close ties with DCDP using Tweedie’s formula in the pixel space. Their algorithm, Diffusive Denoising of Gradient-based Minimization (DDGM), is designed for solving linear tomographic reconstruction problems, which iteratively applies data consistency steps and image denoising through a pre-trained diffusion model. However, they only investigate the use of Tweedie’s formula for regularizing the images, which often results in blurry images (see Section 4.2). Our algorithm circumvents this issue via DDIM which performs diffusion purification with multiple diffusion reverse sampling steps. Furthermore, our framework can incorporate accelerated samplers, latent diffusion models, and consistency models. Consequently, it addresses a broader spectrum beyond tomographic reconstruction and can be readily adapted to a wide range of applications, unlike DDGM.
Hard Data Consistency Methods. Another class of approaches including DiffPIR [33] and ReSample [19] improves upon DPS by alternating between two key steps: (i) at each time step $t$, rather than performing a single-step data consistency update as in DPS, DiffPIR and ReSample solve a multi-step subproblem, which is known as the hard data consistency process, conditioned on the current reconstruction $\bm{x}_{t}$ to obtain an estimate $\bm{x}_{0}$ at the initial timestep, and (ii) map $\bm{x}_{0}$ back to the diffusion trajectory at time $t-1$. While these methods may appear conceptually related to the plug-and-play approaches discussed above, hard data consistency methods generate the final reconstruction within a single reverse diffusion process—meaning the data consistency steps of these methods remain tightly coupled with the reverse sampling. In comparison, our approach employs $K$ iterations of forward and reverse diffusion processes (diffusion purification), demonstrating that data consistency steps can be fully decoupled from the reverse sampling steps. As a result, our method is compatible with one-step CMs, which are not supported by DiffPIR or ReSample.
Red-Diff. RED-Diff [45] is an algorithm that combines regularization by denoising [46] with pre-trained diffusion models within a variational Bayesian framework. It involves solving an optimization problem via gradient descent to minimize a data-fitting loss while maximizing the likelihood of the reconstruction under the diffusion-based denoiser prior.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of our method under various settings. First, we introduce the basic experiment setup in Section 4.1. Second, we show our main experimental results in Section 4.2. Finally, we provide an ablation study in Section 4.3 in terms of noise robustness, computational efficiency, and convergence speed.
@@ 4.1 Experimental Setup
We evaluate our method based on the following setups. The specific choices of hyperparameters are provided in Appendix B.
Inverse Problems Used in Evaluation We evaluate our method on both linear and nonlinear inverse problems. For linear inverse problems, we consider the following tasks with no additive measurement noise: (i) image inpainting using a $100\times 100$ box mask; (ii) 4×4$\times$4 × image super-resolution via uniform downsampling; (iii) Gaussian deblurring with a $61\times 61$ Gaussian kernel of standard deviation $\sigma=3.0$; and (iv) motion deblurring using a $61\times 61$ motion blur kernel with intensity $0.5$. For nonlinear inverse problems, we consider nonlinear deblurring following the experimental setup in [14], with additive Gaussian measurement noise with standard deviation $0.05$.
Figure 3: Pixel space diffusion models for solving various linear inverse problems. These results demonstrate the effectiveness of our algorithm. We can consistently recover images with the highest PSNR.
Evaluation Datasets. For linear inverse problems, we conduct experiments on the FFHQ [22] and LSUN-Bedroom [23] datasets, both of which contain 100 images of size $256\times 256\times 3$. For nonlinear inverse problems, we conduct experiments on ImageNet [24] with 100 images of size $256\times 256\times 3$. For experiments on CMs, we only test on the LSUN-Bedroom dataset and linear inverse problems as a proof-of-concept.
Evaluation Metrics. To quantitatively evaluate the performance, we use standard metrics: (i) Peak Signal-to-Noise Ratio (PSNR), (ii) Structural Similarity Index Measure (SSIM) [47] and (iii) Learned Perceptual Image Patch Similarity (LPIPS) [48].
@@
DDNM [13] $0.233\pm 0.06$ $26.00\pm 2.56$ $0.756\pm 0.08$ $0.105\pm 0.06$ $23.42\pm 3.45$ $0.882\pm 0.08$
RED-Diff [45] $0.267\pm 0.05$ $25.78\pm 1.59$ $0.746\pm 0.07$ $0.115\pm 0.08$ $23.22\pm 3.16$ $0.878\pm 0.05$
DiffPIR [50] $\underline{0.218}\pm 0.06$ $\underline{26.34}\pm 2.34$ $\underline{0.782}\pm 0.06$ $0.108\pm 0.07$ $23.47\pm 1.89$ $\underline{0.884}\pm 0.06$
DCDP-DDIM (Ours) $\bm{0.215}\pm 0.05$ $26.10\pm 3.12$ $0.768\pm 0.08$ $\underline{0.098}\pm 0.01$ $20.79\pm 2.66$ $0.880\pm 0.02$
DCDP-Tweedie (Ours) $0.229\pm 0.05$ $\bm{26.92}\pm 3.33$ $\bm{0.800}\pm 0.07$ $\bm{0.094}\pm 0.02$ $\bm{23.66}\pm 3.42$ $\bm{0.906}\pm 0.02$
@@
DCDP-LDM-DDIM (Ours) $\underline{0.244}\pm 0.05$ $\underline{26.52}\pm 3.21$ $\underline{0.778}\pm 0.08$ $\underline{0.215}\pm 0.04$ $\underline{23.75}\pm 3.32$ $\bm{0.831}\pm 0.05$
DCDP-LDM-Tweedie (Ours) $0.259\pm 0.05$ $\bm{26.88}\pm 3.28$ $\bm{0.789}\pm 0.08$ $0.239\pm 0.04$ $\bm{24.22}\pm 3.65$ $\underline{0.828}\pm 0.05$
DCDP-CM (Ours) $0.208\pm 0.05$ $26.17\pm 3.19$ $0.774\pm 0.08$ $0.108\pm 0.02$ $19.66\pm 2.48$ $0.874\pm 0.02$
Table 1: Quantitative results of super resolution and inpainting on the LSUN-Bedrooms dataset for both pixel space diffusion models and LDMs. The results colored in blue and in orange are for LDMs and CMs respectively, while the others are for pixel space diffusion models. The best results are in bold and second best results are underlined for pixel space models and LDMs, respectively.
Comparison Baselines. We conduct experiments using both pixel space and latent space diffusion models. For the pixel space, we compare our algorithm with DPS [14], Manifold Constrained Gradients (MCG) [11], Plug-and-Play ADMM [49] (PnP-ADMM), DiffPIR [50], RED-Diff [45] and DDNM [13]. For the latent space, we consider Latent-DPS [19], ReSample [19], and Posterior Sampling with Latent Diffusion Models (PSLD) [18]. To ensure a fair comparison, both the baselines and DCDP (ours) use the same sets of pre-trained diffusion models, including pixel space diffusion models from [14] and [7], LDMs from [28], and CMs from [17]. For both DCDP and the baselines considered, we tune the hyperparameters via grid search to obtain the best possible performance.
@@
DDNM [13] $0.217\pm 0.07$ $\underline{27.30}\pm 1.67$ $\underline{0.815}\pm 0.08$
RED-Diff [45] $0.228\pm 0.12$ $23.98\pm 1.23$ $0.687\pm 0.067$ $0.178\pm 0.06$ $24.15\pm 2.34$ $0.704\pm 0.12$
DiffPIR [50] $0.214\pm 0.013$ $27.04\pm 2.67$ $0.801\pm 0.12$ $0.114\pm 0.12$ $27.45\pm 1.89$ $0.866\pm 0.65$
DCDP-DDIM (Ours) $\bm{0.196}\pm 0.05$ $27.13\pm 3.26$ $0.804\pm 0.07$ $\underline{0.065}\pm 0.01$ $\underline{33.56}\pm 3.92$ $\underline{0.947}\pm 0.02$
DCDP-Tweedie (Ours) $\underline{0.212}\pm 0.05$ $\bm{27.74}\pm 3.34$ $\bm{0.825}\pm 0.07$ $\bm{0.067}\pm 0.02$ $\bm{34.47}\pm 4.07$ $\bm{0.956}\pm 0.02$
@@
DCDP-LDM-DDIM (Ours) $0.246\pm 0.05$ $\underline{26.08}\pm 3.02$ $\underline{0.766}\pm 0.07$ $\underline{0.145}\pm 0.03$ $\underline{28.87}\pm 3.77$ $\underline{0.856}\pm 0.06$
DCDP-LDM-Tweedie (Ours) $\bm{0.217}\pm 0.05$ $\bm{27.11}\pm 3.26$ $\bm{0.807}\pm 0.07$ $\bm{0.121}\pm 0.03$ $\bm{29.40}\pm 3.81$ $\bm{0.875}\pm 0.05$
DCDP-CMs (Ours) $0.202\pm 0.05$ $26.91\pm 3.11$ $0.805\pm 0.07$ $0.061\pm 0.02$ $34.10\pm 3.80$ $0.953\pm 0.02$
Table 2: Quantitative results of Gaussian and motion deblurring on the LSUN-Bedrooms dataset for both pixel space diffusion models and LDMs. The results colored in blue and orange are for LDMs and CMs respectively, while the others are for pixel space diffusion models. The best results are in bold and second best results are underlined for pixel space models and LDMs respectively.
4.2 Experimental Results
In this section we present both quantitative and qualitative results for our proposed algorithm and the baselines. For conciseness, we defer the quantitative results for FFHQ to Appendix D. As discussed in Section 3.3, the diffusion purification can be performed by either (i) employing accelerated samplers such as DDIM or (ii) applying a single-step approximation using Tweedie’s formula. We denote these variants as DCDP-DDIM and DCDP-Tweedie, respectively—and as DCDP-LDM-DDIM and DCDP-LDM-Tweedie when implemented with LDMs—and evaluate both of them in our experiments.
Figure 4: Demonstrating the differences between the DCDP methods on SR (4×4$\times$4 ×). The two columns show the reconstructions with PSNRs and their corresponding canny edges.